a rectangular Persian carpet has a perimeter of 240 inches. the length of the carpet is 22 inches more than the width. what are the dimensions of the carpet?
Answer by mananth(16946)

let the width be x
length = x+22
..
Perimeter = 2*L+2*W
2x+2(x+22)=240
2x+2x+44=240
4x+44=240
4x=196
x=49 inches which is the width
length = 49+22=71 inches